OpenStudy (anonymous):

How do I differentiate log x base 2?

OpenStudy (anonymous):

\[\frac{d}{dx}[\log_2(x)]\]?

OpenStudy (campbell_st):

use change of base

OpenStudy (anonymous):

Yes Satellite

OpenStudy (anonymous):

Get lnx/ln2 then ...?

OpenStudy (anonymous):

you can memorize the answer, or write \[\log_2(x)=\frac{\ln(x)}{\ln(2)}\] and differentiate that one don't forget \(\ln(2)\) is just a number

OpenStudy (anonymous):

So does it come down to lnx/xln2 ?

OpenStudy (anonymous):

no

OpenStudy (anonymous):

the derivative of \(\ln(x)\) is \(\frac{1}{x}\)

OpenStudy (anonymous):

therefore you numerator should be \(1\)

OpenStudy (anonymous):

One moment please

OpenStudy (anonymous):

So is it 1/xln2 ?

OpenStudy (anonymous):

yes

OpenStudy (anonymous):

Cheers, three of us discussing which method was right :P